### v2

# SimpleDeepSearcher: Deep Information Seeking via Web-Powered Reasoning Trajectory Synthesis

Abstract
Retrieval-augmented generation (RAG) systems have advanced large language models (LLMs) in complex deep search scenarios requiring multi-step reasoning and iterative information retrieval. However, existing approaches face critical limitations that lack high-quality training trajectories or suffer from the distributional mismatches in simulated environments and prohibitive computational costs for real-world deployment. This paper introduces SimpleDeepSearcher, a lightweight yet effective framework that bridges this gap through strategic data engineering rather than complex training paradigms. Our approach synthesizes high-quality training data by simulating realistic user interactions in live web search environments, coupled with a multi-criteria curation strategy that optimizes the diversity and quality of input and output side. Experiments on five benchmarks across diverse domains demonstrate that SFT on only 871 curated samples yields significant improvements over RL-based baselines. Our work establishes SFT as a viable pathway by systematically addressing the data-scarce bottleneck, offering practical insights for efficient deep search systems. Our code is available at https://github.com/RUCAIBox/SimpleDeepSearcher.
SimpleDeepSearcher: Deep Information Seeking via Web-Powered Reasoning Trajectory Synthesis
Shuang Sun1thanks:    Equal contributions., Huatong Song [^1] footnotemark: 1, Yuhao Wang1, Ruiyang Ren1, Jinhao Jiang1, Junjie Zhang1, Fei Bai1, Jia Deng1, Wayne Xin Zhao1thanks:    Corresponding author., Zheng Liu2, Lei Fang3, Zhongyuan Wang2, Ji-Rong Wen1 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence 3DataCanvas Alaya NeW {sunshuanguns, batmanfly}@gmail.com
1 Introduction
Figure 1: Overall framework of our proposed SimpleDeepSearcher approach. $p$ denotes the prompt, $r$ denotes the reasoning content, $q$ represents the search query, and $d$ refers to the retrieved document after summarization. $t_{s}$ and $t_{e}$ are special tokens indicating the beginning and end of the search query, and $a$ denotes the final answer.
In recent years, retrieval-augmented generation (RAG) methods have significantly enhanced LLMs by incorporating external knowledge retrieval mechanisms (Zhao et al., 2024; Gao et al., 2024). Recent advancements have extended these capabilities to complex deep search scenarios that demand multi-step reasoning with iterative information retrieval and synthesis(Alzubi et al., 2025).
To address the complex reasoning demands in deep search scenarios, early research explored prompt-based strategies that guide models to decompose questions, generate queries, and retrieve information iteratively (Jiang et al., 2024; Teng et al., 2025; Li et al., 2025a). Other studies have attempted to improve model performance through supervised fine-tuning (SFT) (Wang et al., 2025), but due to the lack of high-quality interaction trajectories for reasoning and retrieval, SFT methods often fail to fully activate the model’s autonomous search abilities (Schick et al., 2023). To further enhance the model’s autonomous search capabilities, reinforcement learning (RL) (Sutton et al., 1999) is considered as a promising solution to train models through real-time interaction with the environment (Song et al., 2025; Jin et al., 2025; Zheng et al., 2025). However, most RL-based approaches operate within artificial environments using static document corpora, creating a distributional mismatch with real-world web dynamics. Moreover, the inherent computational intensity of RL training escalates exponentially when interfacing with live search APIs (Sun et al., 2025).
Given the overhead and complexity of RL-based training, we posits that SFT remains a viable pathway for building efficient deep search systems. While SFT offers a streamlined training process, it faces the critical challenge of lacking high-quality training data in deep search scenarios. On the input side, existing QA datasets often lack the diversity and complexity of questions and search-oriented purposes on the Web, which are essential for deep search training. On the output side, traditional answer annotations omit the critical reasoning traces (search operations, evidence synthesis, and efficient decision paths) required for teaching search-integrated reasoning strategies.
In this paper, we propose SimpleDeepSearcher, an efficient search-with-think framework that utilizes strategic data engineering rather than complex training paradigms. Our core methodology centers on a three-fold process for constructing high-quality training data. First, we develop a data synthesis framework grounded in real web search environments, simulating realistic user search behaviors to generate multi-turn reasoning trajectories. Second, we propose a diversity-aware query sampling strategy to optimize domain coverage, semantic complexity, and knowledge unit density. Moreover, we adopt a four-dimensional response curation that enforces format standardization, reasoning efficiency, question difficulty, and search effectiveness. By systematically addressing both query and response-side quality through automated pipelines, SimpleDeepSearcher can obtain high-quality supervised signals based on real web search for complex reasoning to facilitate the SFT process.
Experimental results show that our SFT method significantly boosts model performance on five representative benchmarks with only 871 high-quality training samples. Compared to prompt-based methods, SimpleDeepSearcher achieves a 48.3% improvement, and compared to RL-based RAG methods achieves a 24.9% improvement. This demonstrates that our framework effectively balances performance and efficiency, providing a simple yet powerful approach to enhancing deep search capabilities. Furthermore, our framework is highly extensible that can be combined with other types of training data, the framework is also applicable to RL-based training.
Our main contributions are as follows:
$\bullet$ We propose a real web-based data synthesis framework that simulates realistic user search behaviors, generating multi-turn reasoning and search trajectories.
$\bullet$ We design a multi-criteria data curation strategy that jointly optimizes both input question selection and output response filtering through orthogonal filtering dimensions.
$\bullet$ Experimental results demonstrate that SFT on only 871 samples enables SimpleDeepSearcher to outperform strong baselines (especially RL-based baselines) on both in-domain and out-of-domain benchmarks.
2 Method
2.1 Overview
In this section, we propose SimpleDeepSearcher for complex deep search tasks by leveraging multi-stage data construction strategies.
To address the resource-intensive limitations of deep search systems, we propose SimpleDeepSearcher, a framework that achieves intelligence search through efficient supervised fine-tuning (SFT) with minimal training data. For constructing high-quality SFT data, we establish a systematically designed data synthesis and curation pipeline, as illustrated in Figure 1.
First, we replace static document retrieval with real-time network interactions, simulating human search behavior through an iterative cycle of “reasoning-searching-summarizing-generating.” By directly processing raw HTML content via commercial search APIs, we capture diverse web information features—ranging from structured data snippets to unstructured narrative discourse. Based on this, we first filter input queries using domain heterogeneity, keyword diversity, and knowledge unit complexity to construct a maximally informative training foundation while ensuring selected queries align with real-world web search scenarios. Additionally, we apply a filtering mechanism to LLM-synthesized responses, implementing a four-dimensional quality filter that simultaneously optimizes format standardization, reasoning path control, question difficulty, and search effectiveness to guarantee response quality.
The framework’s modular design offers three distinctive advantages over conventional approaches: First, it exposes the model to authentic search artifacts and noise patterns through real web interactions. Second, our multidimensional filtering strategy enables state-of-the-art performance with remarkably small SFT datasets, eliminating dependency on resource-heavy RL training. Third, the decoupled architecture between data synthesis and model constraints provides exceptional flexibility that our curated datasets can enhance any LLMs while maintaining compatibility with emerging reasoning architectures and alternative training paradigms including RL. Since the searched content is not generated by the LLM, we mask out these tokens during the SFT process.
Our methodology achieves unprecedented efficiency in search-oriented model training, reducing computational demands while maintaining competitive performance through strategic data quality optimization rather than brute-force data quantity.
2.2 Data Synthesis in Real Web Environment
Typically, traditional retrieval-augmented generation (RAG) systems rely on closed and static knowledge corpora. Such knowledge corpora exhibit two primary limitations: firstly, the content they contain often consists of refined and condensed segments (Barnett et al., 2024); secondly, the information within these knowledge corpora lacks timeliness (Ouyang et al., 2025). Consequently, RAG systems are limited in their ability to simulate authentic user search behaviors, as users typically search within open, dynamic, and complex web environments where the information is not only diverse in format and varied in quality but is also frequently accompanied by redundancy and noise. In light of this, our data synthesis approach does not rely on curated document collections but is instead grounded in the real, open web environment. This authentic web environment also places greater demands on the model’s capabilities for information extraction, synthesis, and reasoning.
Building upon the widely adopted iterative deep search process (Li et al., 2025a) of reason-search-summarize-generate, we develop an automated pipeline for large-scale training data synthesis. For each query, our framework at each iteration (1) initiates web searches via commercial APIs, (2) extracts and processes raw HTML content, (3) applies an LLM to reason over multi-source evidence, and (4) continues for the next iteration or stop iteration. By sampling multiple reasoning paths per query, we capture nuanced decision-making processes inherent to real-world information synthesis.
Our data synthesis strategy is firmly rooted in real web scenarios, which substantially enriches the diversity and authenticity of training samples. Our strategy generates more practical and representative supervisory signals for SFT, thereby addressing critical limitations in conventional SFT paradigms.
2.3 Diversity-aware Query Sampling
Algorithm 1 Diversity-aware Query Sampling
1:Annotated dataset $D$ with domains $d_{1},d_{2},\dots,d_{m}$, target number of queries $N$
2:$N_{d}\leftarrow N/m$
3:$S\leftarrow\emptyset$triangleright Initialize the target set
4:for $i$ = 1 to $n$ do
5:     $D_{d_{i}}\leftarrow\{x\in D\mid\text{domain}(x)=d_{i}\}$
6:     Sort $D_{d_{i}}$ by descending interrogative
7:     words
8:     while $|S_{d_{i}}|<N_{d}$ and $D_{d_{i}}\neq\emptyset$ do
9:         $K\leftarrow\emptyset$triangleright Initialize the keyword set
10:         for each sample $x$ in $D_{d_{i}}$ do
11:              if $|S_{d_{i}}|\geq N_{d}$ then
12:                  break
13:              end if
14:              $kw\leftarrow\text{keywords}(x)$
15:              if $x\notin S$ and $kw\cap K=\emptyset$ then
16:                  $S\leftarrow S\cup\{x\}$
17:                  $K\leftarrow K\cup\text{keywords}(x)$
18:                  $D_{d_{i}}\leftarrow D_{d_{i}}\setminus\{x\}$
19:              end if
20:         end for
21:     end while
22:end for
23:return $S$
To engineer a deep search architecture with advanced query comprehension and reasoning capabilities, we implement a strategic repurposing of open-domain question answering (QA) resources. These curated datasets offer natural language questions that inherently require multi-hop information retrieval operations, thereby exhibiting strong task alignment with the cognitive demands of deep search systems. Our selection protocol combines single-hop and multi-hop QA benchmarks through principled composition, ensuring coverage of both atomic and composite reasoning paradigms.
However, empirical evidence suggests that naive dataset scaling yields diminishing returns in SFT. The efficacy of such approaches fundamentally depends on the intrinsic diversity and informational entropy of training instances. While existing open-domain QA corpora provide substantial volume, systematic analysis reveals three critical limitations: (1) domain-specific overrepresentation creating skewed knowledge distributions, (2) repetitive syntactic patterns reducing linguistic variability, and (3) semantic simplicity thresholds below real-world query complexity. These factors collectively induce model brittleness and constrain cross-domain generalization potential. To address these critical limitations, we introduce a diversity-aware query sampling strategy to implement systematic data filtering through tripartite orthogonal criteria:
Domain Heterogeneity encompasses the systematic classification of query semantics across distinct knowledge domains (e.g., history, science, politics). This dimension ensures a balanced distribution of questions across different domains, thereby reducing domain-specific biases and enhancing generalization capabilities.
Keyword Diversity focuses on the distributional diversity of core semantic constituents. we ensure non-redundant exposure to low-frequency conceptual entities, multi-order relational dependencies, and contextually ambiguous referential expressions. Such systematic variation compels the model to transcend superficial lexical pattern matching, instead developing reasoning architectures essential for interpreting complex entity interactions.
Complexity of knowledge units captures the frequency of interrogative terms used in questions (e.g., what, when), which serve as indicators of syntactic and semantic complexity. Questions with greater inquiry potential are given priority, ensuring comprehensive modeling of implicit reasoning chains triggered by diverse question formulations.
We developed a systematic query selection framework incorporating three complementary dimensions: domain heterogeneity, keyword diversity, and complexity of knowledge units. First, we partition the dataset into domain-specific clusters using the LLM-generated semantic classifications. Within each domain cluster, queries are ranked by knowledge unit complexity scores derived from conceptual density analysis. Subsequently, we perform iterative selection using a greedy algorithm that maximizes keyword diversity while maintaining inter-domain balance. The detailed procedure for query sampling is presented in Algorithm 1.
2.4 Multi-Dimention Response Curation
Building upon the aforementioned data synthesis and query sampling strategies, we have successfully generated high-quality training data derived from real-world web environments. However, due to the inherent unpredictability of LLM reasoning, the quality of synthesized data exhibits considerable variability despite meticulous control over input and generation processes. Three primary issues are observed: (i) Formatting irregularities, such as inconsistent reasoning languages, non-standard formats for search and reasoning steps, and heterogeneous answer formats; (ii) reasoning redundancy, including hypothesis overgeneration, fabricated retrieval content, and excessive validation loops; (iii) inefficient search strategies, including redundant search exploration, contextual myopia and failure to retrieve relevant information.
The presence of low-quality reasoning outputs in language models not only compromises performance and transparency but also introduces noise into training signals, leading to inefficient computational resource utilization. To address these challenges, we developed a systematic filtering protocol that selects optimal solutions through rigorous and comprehensive evaluation of multiple responses per query.
To mitigate these issues, we impose strict constraints on both the format and content of sampled responses, retaining only those that satisfy all predefined criteria. Our filtering strategy, structured around four pillars, ensures retention of high-quality reasoning data while promoting efficient search integration.
Format Standardization. Filter out responses with mixed reasoning languages or incorrect reasoning and search formats, and correct answers with formatting errors to ensure consistency and standardization across all responses. Responses exhibiting mixed languages, irregular reasoning structures, or formatting inconsistencies were excluded. Automated correction aligned remaining answers with standardized templates.
Reasoning Path Control. Strictly limit the use of reflection expressions (e.g., alternatively, wait, etc.) and control the length of reasoning to avoid unnecessary and redundant reasoning steps. Reasoning models tend to hypothesize, infer, and reflect based on internal knowledge, often resulting in delayed use of search tools and inefficient reasoning. By regulating the reasoning path, the model can learn to seamlessly integrate search into its inference process and adopt more efficient reasoning strategies.
Question Difficulty. Filter out questions with consistently high accuracy across multiple reasoning attempts and prioritize those with lower accuracy. Accuracy obtained from multiple samples can serve as a proxy for question difficulty. Selecting more challenging questions helps enhance the model’s ability to handle complex queries.
Search Effectiveness. Among multiple candidate responses, prioritize those with fewer search steps and more diverse search content. This encourages the model to not only invoke search capabilities but also to learn how to formulate effective sub-queries based on the original question for efficient information retrieval.
Based on the above dimensions, we first collect metadata for each response, such as the number of search steps, reasoning length, and accuracy. Subsequently, responses are filtered sequentially based on format standardization and reasoning path control. Then, based on question difficulty, questions with high accuracy are removed. For each remaining question, we retain multiple high-quality responses that meet all constraints and sort them by search steps. According to search effectiveness, the response with the fewest search steps is selected as the final answer. Through this process, we ultimately obtained 871 high-quality question-answer pairs. This multi-criteria approach not only enhances model training efficiency but also provides insights into optimal human-AI reasoning patterns.
Models Methods 2Wiki MuSiQue Bamboogle Frames GAIA
F1 LasJ F1 LasJ F1 LasJ F1 LasJ F1 LasJ
Qwen-7B Directly Gen 27.7 26.8 9.6 6.2 18.2 17.6 12.6 10.1 13.6 6.8
Standard RAG 34.8 34.8 17.2 14.6 31.5 31.2 13.9 13.5 - -
Search-o1 48.0 51.2 21.5 20.6 57.9 59.2 30.9 35.0 24.3 21.4
R1-Searcher 63.4 66.4 29.0 26.8 68.2 68.8 34.4 40.3 24.1 20.4
DeepResearcher 59.7 66.6 27.1 29.3 71.0 72.8 - - - -
SimpleDeepSearcher 70.6 79.8 28.2 29.4 74.5 76.8 44.9 55.3 39.3 36.9
Qwen-32B Directly Gen 31.7 31.2 13.3 12.4 25.7 25.6 15.6 14.2 18.6 13.9
Standard RAG 43.7 45.0 19.5 16.8 40.8 40.8 19.4 19.4 - -
Search-o1 64.9 74.8 29.1 30.6 74.4 78.4 47.2 56.8 36.5 34.0
SimpleDeepSearcher 71.9 81.2 30.6 33.0 78.1 80.0 50.1 60.8 42.1 40.8
DDQ-32B Directly Gen 36.9 36.2 19.6 16.0 32.6 32.8 27.8 29.2 14.8 9.7
Standard RAG 48.1 50.0 24.0 21.6 42.6 46.4 26.5 28.9 - -
Search-o1 49.6 55.2 25.4 23.8 65.7 68.0 32.2 38.7 23.2 24.3
SimpleDeepSearcher 69.0 77.4 32.9 33.6 80.5 83.2 52.2 63.8 42.0 41.7
QwQ-32B Directly Gen 39.6 39.8 18.9 17.4 29.6 29.6 28.1 31.3 16.8 11.7
Standard RAG 48.4 50.6 21.8 19.4 42.5 46.4 27.4 31.6 - -
Search-o1 69.4 78.0 34.3 36.4 78.7 78.4 51.6 64.4 38.3 37.9
SimpleDeepSearcher 75.6 84.4 34.8 37.4 83.4 88.0 56.8 68.8 48.9 50.5
Table 1: Performance comparisons between SimpleDeepSearcher and the baselines on QA benchmarks. The best results are in bold and the second-best are underlined. ${}^{\dagger}/\ddagger$ represents in-domain/out-domain datasets. Results marked with * are cited from their official paper or report. Qwen-7B, Qwen-32B, DDQ-32B are the abbreviations of Qwen-2.5-7B-Instruct, Qwen-2.5-32B-Instruct, and Deepseek-Distilled-Qwen-2.5-32B, respectively.
3 Experiments
3.1 Experimental Setup
Datasets.
We sample training data from single-hop and multi-hop knowledge-intensive QA datasets to cover a wide range of domains and question difficulty. For single-hop questions, we use Natural Questions (Kwiatkowski et al., 2019) and SimpleQA (Wei et al., 2024). For multi-hop questions, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), MuSiQue (Tang and Yang, 2024), and MultiHopRAG (Tang and Yang, 2024). To test the model’s performance on out-of-domain data, we select Bamboogle (Press et al., 2022), FRAMES (Krishna et al., 2024), and GAIA (Mialon et al., 2023) as extra test sets. These datasets are not used during training and help evaluate how well the model works on new domains. We evaluate our approach on 500 randomly sampled instances from the validation sets of HotpotQA, 2WikiMultiHopQA, and MuSiQue. For Bamboogle and FRAMES, we use their full test sets. For GAIA, we use 103 examples from the text-only validation subset (Li et al., 2025b).
Metrics.
We report results using two metrics: F1 score and LLM-as-Judge (LasJ). The F1 score captures the word-level similarity between the predicted and golden answers, while LasJ leverages GPT-4o-mini to evaluate the correctness of the predicted response.
Baselines.
We consider following type of baselines: Naive Generation: Direct generation of answers without retrieval. Standard RAG (Zhao et al., 2024): Directly retrieve relevant documents by querying the original question. Search-o1 (Li et al., 2025a): Encourages the model to perform self-initiated retrieval using prompts. RAG-RL: R1-Searcher (Song et al., 2025) and DeepResearcher (Zheng et al., 2025), the open-source 7B model trained with reinforcement learning to enable self-initiated retrieval. We conduct experiments using the following model backbones with an online search engine, including Qwen-2.5-7B-Instruct, Qwen-2.5-32B-Instruct, Deepseek-Distilled-Qwen-2.5-32B, and QwQ-32B.
Implementation Details.
Our experimental setup consists of four main components: query sampling, data synthesis, generation, and SFT. During query sampling, we used QwQ-32B to annotate each query with its corresponding domain and keywords. For data synthesis, we employed QwQ-32B as the reasoning model and Google Search API as the search engine, with a maximum of 10 search calls and 15 reasoning turns per query. For each query, we sampled 10 candidate responses. For generation, all models are configured with a maximum sequence length of 20,480 tokens, temperature of 0.6, top-p of 0.95, and top-k of 40. In the SFT phase, we use a total batch size of 64 and train for 6 epochs with a learning rate of 1e-5, warmup ratio of 0.03, and a sequence length of 30,000 tokens. During fine-tuning, external retrieval documents are masked to avoid learning from noisy or spurious information. All prompts used in the experiments are provided in Appendix E.
3.2 Main Results
Table 1 presents the main results of the proposed SimpleDeepSearcher and baselines across five representative datasets.
Firstly, SimpleDeepSearcher consistently outperforms all existing baseline methods across five benchmark datasets. Specifically, it achieves the best performance not only on in-domain datasets (i.e., 2Wiki, MuSiQue) but also shows substantial improvements on out-of-domain datasets (i.e., Bamboogle, FRAMES, GAIA), demonstrating its strong generalization ability.
Besides, SimpleDeepSearcher consistently outperforms reinforcement learning-based methods such as R1-Searcher and DeepResearcher across most evaluation metrics. These approaches are trained on large-scale datasets using complex reinforcement learning algorithms. In contrast, our method relies on supervised fine-tuning with only 871 training examples. This demonstrates that our framework achieves strong performance while maintaining high data efficiency, offering a simple yet effective alternative for improving deep search capabilities.
Category Method Bamboogle GAIA
F1 LasJ F1 LasJ
Ours 74.5 76.8 39.3 36.9
Query Sampling w/o DH 69.7 70.4 35.6 35.8
w/o KD 73.2 76.0 32.9 31.1
w/o CKU 71.7 74.4 32.1 29.1
Environment w/o Online 74.0 74.4 30.4 28.2
Response Curation w/o FS 72.8 75.2 38.0 36.9
w/o RPC 71.7 74.4 31.6 30.1
w/o QD 67.1 70.4 32.9 32.0
w/o SE 72.6 73.6 37.7 35.0
Table 2: Results of variants of SimpleDeepSearcher on Bamboogle and GAIA.
Thirdly, SimpleDeepSearcher achieves stable and substantial performance improvements across models with diverse backbones and parameter scales, ranging from 7B to 32B. For instance, compared to Search-o1, it achieves relative improvements of 48.3%, 42.6%, and 11.5% on Qwen2.5-7B-Instruct, DeepSeek-R1-Distill-Qwen-2.5-32B, and QwQ-32B, respectively. This demonstrates the strong generalization ability of our distillation and self-distillation strategies, with the selected data consistently leading to performance gains across heterogeneous model architectures.
4 Further Analysis
4.1 Ablation Study
To validate the effectiveness of the proposed SimpleDeepSearcher, we conduct a comprehensive ablation analysis using Qwen2.5-7B-Instruct on the Bamboogle and GAIA datasets. We conduct detailed ablation studies on three main aspects: (1) Query Sampling: w/o DH removes domain heterogeneity filter, w/o KD removes keyword diversity filter, w/o CKU removes complexity of knowledge units filter; (2) Environment: w/o Online uses local dense dense retrieval to synthesize training data; (3) Response Curation: w/o FR removes format regularization filter, w/o RPC removes reasoning path control filter, w/o QD removes question difficulty filter, w/o SC search count filter. As shown in Table 2, all ablated variants exhibit a decline in performance compared to our full method, underscoring the integral contribution of each component. Among them, w/o QD leads to the most significant performance drop, suggesting that question difficulty plays a crucial role in training. More challenging questions are more likely to stimulate the model’s autonomous retrieval capabilities during reasoning.
Method Bamboogle GAIA
F1 LasJ F1 LasJ
Distilled (Ours) 74.5 76.8 39.3 36.9
w. DPO 75.0 79.2 39.0 37.9
w. Reinforce++ 73.8 75.8 29.4 24.3
Table 3: Evaluation Results of RL-based Methods.
Model #Alternatively #Search Output Length
QwQ-32B 7.933 2.390 867.148
QwQ-32B-SFT 4.051 2.329 581.731
Table 4: Statistical analysis of model outputs.
4.2 Effect of Post-SFT RL
Recent studies have investigated the integration of RL and RAG (Song et al., 2025; Jin et al., 2025; Zheng et al., 2025). We further examine the advantages and limitations of applying RL after SFT.
We apply DPO and REINFORCE++ to conduct offline and online reinforcement learning, respectively. As shown in Table 3, the model trained with DPO achieves further improvements over the SFT baseline, demonstrating the effectiveness of offline preference optimization. In contrast, the model trained with REINFORCE++ produces significantly shorter responses (see Appendix C) and shows notable performance degradation on both the Bamboogle and GAIA benchmarks. This suggests that online RL mainly triggers retrieval behavior, but brings little benefit to models that are already good at retrieval. We hypothesize that the success of offline DPO stems from its ability to leverage high-quality trajectories generated by a strong LLM. These trajectories provide informative preference signals and stable supervision, allowing the model to refine its reasoning and search strategies.
Model Plan. Search Summ.
Qwen-7B 0.416 0.455 0.363
Qwen-7B-SFT 0.590 0.677 0.584
QwQ-32B 0.623 0.680 0.594
QwQ-32B-SFT 0.629 0.713 0.624
Table 5: Proportion of instances containing the correct answer at each stage of the inference process (Planning, Search, and Summarization), before and after SFT.
4.3 Response Redundancy
In this part, we analyze how SFT impacts redundant reasoning and search behavior. Specifically, we focus on three indicators: (1) the frequency of the reflective word “alternatively”, which signals hesitation or divergent reasoning; (2) the average length of reasoning chains, measured by output length; and (3) the number of search calls made during inference. Our analysis is based on the QwQ-32B model, evaluated on the 2Wiki, MuSiQue, and Bamboogle datasets. As shown in Table 4, the average use of “alternatively” and the overall output length are both significantly reduced after SFT. Moreover, the model issues fewer search queries. These results indicate that our self-distillation approach improves both the reasoning clarity and search efficiency of the model. This improvement can be attributed to the high-quality training data selected through our proposed method.
4.4 Stage-wise Analysis
In this part, we analyze how training improves the performance of each sub-task in our approach, including iterative search, planning, and summarization. We evaluate the proportion of cases in which the final answer appears during each sub-process to quantify the efficiency of that stage. To eliminate interference from the summarization stage, all summarization models are kept identical during inference, with detailed settings provided in Appendix D. The results are shown in Table  5. We can observe substantial improvements across all components, with the search component showing the most significant gain. This suggests that training effectively enhances the model’s ability to generate more coherent reasoning and search trajectories, leading to more accurate information retrieval and improved overall model performance.
Models Summarization Models Bamboogle GAIA
F1 LasJ F1 LasJ
Qwen-7B-SFT before training 70.8 71.2 28.0 26.2
after training 67.5 68.8 23.9 21.4
QwQ-32B 74.5 76.8 39.3 36.9
GPT-4o-mini 70.9 76.8 33.7 32.0
QwQ-32B-SFT before training 83.5 88.0 48.9 50.5
after training 83.9 86.4 43.2 47.6
GPT-4o-mini 80.0 80.8 40.5 44.7
Table 6: Performance comparison across two benchmarks using different summarization models.
Training Data Bamboogle GAIA AIME
F1 LasJ F1 LasJ F1 LasJ
- Reasoning 74.5 76.8 39.3 36.9 13.3 13.3
+ Reasoning 76.9 80.8 37.2 37.9 20.0 20.0
Table 7: Results of the SimpleDeepSearcher trained w/ and w/o reasoning data across three benchmarks.
4.5 Effect of Summarization Model
This part investigates the impact of the summarization model on overall performance. We fix the reasoning model and conduct a comparative analysis of overall performance using different summarization models. As shown in Table 6, QwQ-32B demonstrates the strongest summarization capability and is therefore selected as the summarization model for all reasoning models. Furthermore, using fine-tuned models for summarization leads to performance degradation on downstream tasks compared to their pre-trained counterparts. This might be attributed to the reduced long-text summarization ability of the fine-tuned models, due to the distributional shifts on a limited task and domain of the training data. This decline is more pronounced for models with fewer parameters.
4.6 Effect of Additional Reasoning Data
We further investigate the impact of incorporating complex mathematical reasoning data on Qwen2.5-7B-Instruct. As shown in Table 7, this leads to consistent performance gains across all benchmarks. Furthermore, Figure 2 and Table 8 reveals significant alterations in the model’s behavioral patterns on two kinds of tasks: for tasks emphasizing complex reasoning (e.g., AIME, GAIA), the model generates longer and more in-depth reasoning outputs; for search tasks (e.g., Bamboogle), the model performs more searches and explores more thoroughly. These findings suggest that incorporating complex reasoning data helps the model learn to adapt its reasoning and search strategies to the specific demands of a task. This adaptability is critical for addressing complex and diverse queries.
Figure 2: Average reasoning length across three benchmarks w/ and w/o reasoning data for training.
Training Data Search Count
Bamboogle GAIA AIME
- Reasoning 1.552 1.757 0
+ Reasoning 1.672 1.845 0
Table 8: Average search count across three benchmarks of the model trained w/ and w/o reasoning data.
5 Conclusion
In this work, we present SimpleDeepSearcher, a lightweight yet effective framework for deepsearch tasks, addressing the limitations of existing RAG methods that rely heavily on complex training paradigms or suffer from distributional mismatches. By leveraging realistic web search simulations and a multi-criteria data curation strategy, we construct high-quality training trajectories that enable efficient supervised fine-tuning. Despite using only 871 curated samples, our method achieves substantial gains over RL-based baselines across diverse in-domain and out-of-domain benchmarks. Our results highlight the potential of strategic data engineering to empower deep search reasoning.
Limitation
Despite our substantial efforts, this work is subject to two limitations stemming. Due to limitations in training resources and hardware, we conducted distillation training on 7B and 32B models. In future work, we plan to train and evaluate our framework on larger-scale models (i.e., 72B) to further verify its generalization capability and robustness. Additionally, because of the inherent difficulty in synthesizing multi-hop data, the original data used for distillation primarily consisted of relatively simple multi-hop questions. If more realistic and challenging multi-hop queries can be synthesized in the future, applying our framework for filtering and training may yield even better performance.
Appendix A Related Work
Retrieval-Augmented LLMs. To improve the factual precision of LLM-generated texts (Zhao et al., 2024), researchers enhance LLMs with retrieval-augmented generation (RAG) (Guu et al., 2020). Various approaches have been proposed, such as branching-based methods (Kim et al., 2024), summarization-based methods (Li et al., 2023), and adaptive retrieval techniques (Jeong et al., 2024). With the increase in model parameters, LLMs have demonstrated chain-of-thought reasoning capabilities Wei et al. (2022), and many researchers to integrated such reasoning with RAG via prompt engineering (Shao et al., 2023; Trivedi et al., 2023). Other studies have attempted to distill retrieval abilities into smaller models through supervised fine-tuning Asai et al. (2024). However, these approaches limit the model’s capacity with a fixed reasoning path.
Enhancing LLMs with Search. Recently, several deep search frameworks are proposed (Ren et al., 2025; Alzubi et al., 2025). They integrate large language models with search engines in a more flexible and dynamic manner. Search-o1 (Li et al., 2025a) simulates deep search in LLMs through prompt engineering, allowing them to retrieve information independently during multi-step reasoning. R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) equip large language models with retrieval tools and train them end-to-end using reinforcement learning. This approach effectively enhances the model’s ability to interleave reasoning with retrieval during inference. However, due to the inherent complexity of RL and its high computational demands, conducting large-scale experiments on full-sized LLMs remains challenging. SimpleDeepSearcher synthesizes high-quality training data via broad query sampling and precise filtering, enabling strong deep search performance with minimal training cost.
Appendix B DPO Detailed Settings
Our objective was to identify answer trajectories that were both correct and demonstrated efficient reasoning and search paths. To this end, we construct preference pairs $(R_{w},R_{l})$, where $R_{w}$ denotes the preferred trajectory and $R_{l}$ the rejected one. We repurpose our previously established pipeline for query sampling and data synthesis. During the data synthesis stage, we generate responses using the strongest SFT-trained model, SimpleDeepSearcher-QwQ-32B-SFT, and the target model to be optimized, SimpleDeepSearcher-Qwen-7B-SFT. Responses generated by SimpleDeepSearcher-QwQ-32B-SFT that pass both the formatting and reasoning path control checks are treated as chosen examples, while those generated by SimpleDeepSearcher-Qwen-7B-SFT that fail these checks are treated as rejected examples. Ultimately, we construct a dataset consisting of approximately 875 training pairs.
For Direct Preference Optimization (DPO) training, we utilize a learning rate of $5\times 10^{-7}$, a $\beta$ of $0.1$, training for $5$ epochs with a batch size of $256$, a warm-up ratio of $0.1$, and a maximum sequence length of $10000$.
Appendix C REINFORCE++ Detailed Settings
Figure 3: Changes in Sequence Length and Reward During REINFORCE++ Training.
To construct the reinforcement learning (RL) dataset, we utilized the model that had been trained though SimpleDeepSearcher to perform rollout sampling on the training sets of 2Wiki and HotpotQA. For each question, eight candidate responses were generated. From this pool, we selected 2480 samples corresponding to questions with one to six correct answers, ensuring diversity in the RL training data.
The reward function employed in REINFORCE++ consists of two components: an answer reward and a format penalty. The answer reward is calculated as the F1 score between the predicted answer and the reference answer, providing a direct measure of response accuracy. In addition, a discrete format penalty of $-2$ is applied if any of the following undesirable behaviors are detected:
$\bullet$ Self-Retrieved Content: The model fabricates content that is not retrieved from external sources.
$\bullet$ Contains Gibberish: The generated output contains nonsensical, irrelevant, or corrupted text segments.
$\bullet$ Excessive Analytical Markers: The response contains more than 5 occurrences of phrases such as Alternatively, Wait, or Hmm, which are treated as signals of incoherent reasoning.
$\bullet$ Lack of Boxed Answers or Excessive Reasoning Length: The model either executes more than 8 retrieval steps or the token length of the analytical content between any two retrievals exceeds 8,096 tokens.
If none of these conditions are met, no penalty is applied. To maintain on-policy training throughout the RL process, we adjusted the batch size to ensure that learning was based on the most recent policy rollouts. Figure 3 shows the variations in response length and reward values observed during the training process.
Appendix D Stage-wise Analysis Settings
We conduct a comparative analysis of Qwe2.5-7B-Instruct and QwQ-32B before and after training across the 2Wiki, MuSiQue, and Bamboogle benchmarks. During inference, we fix the summarization model to QwQ-32B across all comparisons to eliminate potential interference from the summarization component (the impact of the summarization model will be further discussed in Section 4.5).
Appendix E Instruction Templates
Figure 4 shows the instruction for annotating the domain and keywords of questions. Figure 5 shows the instruction for LLM as a judge. Figure 6 shows the instruction for the reasoning model. Figure 7 shows the instruction for the summarization model.
Figure 4: Instruction for annotation.
Figure 5: Instruction for LLM as a judge.
Figure 6: Instruction for reasoning model.
Figure 7: Instruction for summarization model.